OpenStudy (anonymous):

please help with Midterm Q1 The elements in this circuit have the following values: R1=200Ω, R2=200Ω, R3=100Ω, V1=5V, I1=0.001A and I2=0.005A

OpenStudy (anonymous):

a)2.84V b) 0.68V c).02332 d)0.02332 e)0.004624 f)0.0085 g).00068 h)-0.0034

OpenStudy (anonymous):

thank you very much Vivek26.. !!!! but please.. the value f) 0..0085 is not correct here .. do u know why ?

OpenStudy (anonymous):

the current through R1 and R2 is 0.0017A and V1 is 5 power=V*I=0.0017*5=.0085

OpenStudy (anonymous):

if you have any doubt feel free to ask

OpenStudy (anonymous):

What is the power (in Watts) coming out of the voltage source V1? 0.054

OpenStudy (anonymous):

Thank you very much Vivek2609 .. and 1431 mahi !!!